OpenStudy (anonymous):

12-3(2w+1)=7w-3(7+w)

OpenStudy (anonymous):

distribute the -3 into the parenthesis (2w+1)

OpenStudy (anonymous):

then distribute the other -3 into the parenthesis (7+w)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

write what you get.

OpenStudy (anonymous):

-6w+3 and -21+-3w

OpenStudy (anonymous):

great, now put that back with the rest of the equation.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and then combine like terms on either side of the equation

OpenStudy (anonymous):

i got 15-6w and 4w-21

OpenStudy (anonymous):

10w-21 for the second one not 4w-21

OpenStudy (anonymous):

is 4w sorry

OpenStudy (anonymous):

checking

OpenStudy (anonymous):

oops, your answer of -6w+3 was incorrect, it should have been -6w+(-3)

OpenStudy (anonymous):

in the next step you were on the right track. however, \[12-6w+(-3)=7w-10+(-3w)\]

OpenStudy (anonymous):

\[12-3-6w=7w-3w-10\]

OpenStudy (anonymous):

\[9-6w=4w-10\]

OpenStudy (anonymous):

following?

OpenStudy (anonymous):

ok i fix it

OpenStudy (anonymous):

do you know what to do next?

OpenStudy (anonymous):

now for 9-6w=4w-21thats is where i am

OpenStudy (anonymous):

my final result was 3